OpenStudy (anonymous):

The area of a square garden is 200 m². How long is the diagonal?

OpenStudy (anonymous):

@Jack1 @Yttrium

OpenStudy (yttrium):

Solve for it's side and use pythagorean theorem.

OpenStudy (jack1):



OpenStudy (anonymous):

need a little more help i can do the work its just setting it up is the problem

OpenStudy (anonymous):

@Yttrium

OpenStudy (anonymous):

@troyandthor

OpenStudy (anonymous):

heeeeeeelllp

OpenStudy (yttrium):

A = s^2, right? Therefore s = sqrt(A) which is s = sqrt(200) s = 14.142 So, using pythagorean theorem c^2 = 2s^2 (note that s = s since it is a square) Therefore, c^2 = 2(14.142)^2 and c = sqrt (399.992) Hence c = 19.999 Get it?

OpenStudy (anonymous):

Yeah, there ya go^

OpenStudy (yttrium):



OpenStudy (anonymous):

its 20m

OpenStudy (anonymous):

Hey @Yttrium can you help me with my problem? http://openstudy.com/updates/52370875e4b0b9d6b2d666d0

OpenStudy (yttrium):

Yes @imani1993